# Time Travel is Cheating: Going Live with DeepFund for Real-Time Fund Investment Benchmarking

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated notable capabilities across financial tasks, including financial report summarization, earnings call transcript analysis, and asset classification. However, their real-world effectiveness in managing complex fund investment remains inadequately assessed. A fundamental limitation of existing benchmarks for evaluating LLM-driven trading strategies is their reliance on historical back-testing, inadvertently enabling LLMs to “time travel”—leveraging future information embedded in their training corpora, thus resulting in possible information leakage and overly optimistic performance estimates. To address this issue, we introduce DeepFund, a live fund benchmark tool designed to rigorously evaluate LLM in real-time market conditions. Utilizing a multi-agent architecture, DeepFund connects directly with real-time stock market data—specifically data published after each model’s pretraining cutoff—to ensure fair and leakage-free evaluations. Empirical tests on nine flagship LLMs from leading global institutions across multiple investment dimensions—including ticker-level analysis, investment decision-making, portfolio management, and risk control—reveal significant practical challenges. Notably, even cutting-edge models such as DeepSeek-V3 and Claude-3.7-Sonnet incur net trading losses within DeepFund real-time evaluation environment, underscoring the present limitations of LLMs for active fund management. Our code is available at https://github.com/HKUSTDial/DeepFund.
@@ 1 Introduction
The financial industry has witnessed an AI-driven revolution over the past decade [27, 31, 59, 47]. Advanced machine learning techniques, especially Large Language Models (LLMs), have transformed practices ranging from high-frequency trading algorithms [34, 5, 63, 32], risk assessment models [13, 60, 20], to investment decisions [52, 46, 56, 58, 26, 54], fundamentally changing how financial institutions operate and make decisions.
@@
A fundamental limitation in extending these evaluations to trading performance lies in their reliance on retrospective back-testing. Back-testing is the standard method for assessing trading strategies [3, 15], but it becomes problematic when applied to LLM-driven strategies because the model may have been pre-trained on the very historical data used for testing, which leads to severe information leakage issue [55, 42, 39, 11]. Undoubtedly, an LLM can appear to perform extraordinarily well on historical market data simply by regurgitating events it has already seen, rather than genuinely predicting outcomes. In other words, the model can effectively “time travel” by using future knowledge during evaluation—a form of cheating that inflates its apparent performance. This issue is exacerbated by the varying knowledge cut-off dates of different LLMs.
@@
DeepFund: Fund Benchmark Going Live. To address the critical gap identified above, we introduce DeepFund – a comprehensive framework for real-time fund investment benchmarking, as shown in Figure 1(b). Inspired by previous works [33, 19, 50], DeepFund assesses LLM’s ability to make effective investment decisions in a live-market environment, explicitly preventing any leakage of future data. In particular, our approach offers three key contributions:
@@ 2 DeepFund: Multi-Agent Fund Investment Going Live
To facilitate seamless and flexible data ingestion, we implement a modular API gateway that interfaces with multiple financial data providers, such as Yahoo Finance and Alpha Vantage. This modularity (refer to Appendix D.5) ensures adaptability to diverse data sources and straightforward integration into varying market contexts. By offering direct feedback from live market conditions, the environment guarantees authenticity in the decision-making process, fundamentally shifting the evaluation paradigm from retrospective back-testing to dynamic, real-time interaction.
Single Agent Design. Similar to previous works [30, 46, 52, 26], each agent in our framework, powered by the selected LLM backend, fulfills a specific role within the investment process:
@@
Portfolio Manager: Integrates multiple analyst signals to make executive investment decisions (Buy, Sell, Hold), manages risk control, and maintains a dual-memory architecture [41, 44, 23] (see Appendix D.2) to reflect historical transactions and current portfolio states.
@@
Multi-Agent Workflow. The multi-agent workflow [35, 57, 30, 49] adopts an orchestrator-worker paradigm [^1] See Anthropic blog https://www.anthropic.com/engineering/building-effective-agents to mimic a realistic fund management process. Initially, the Financial Planner selects and assigns analysts based on real-time market conditions and portfolio status. Next, the selected Analyst Team concurrently processes domain-specific information and generates structured analytical signals. Finally, the Portfolio Manager synthesizes these signals, evaluates portfolio risks, decides optimal trading actions, and updates the investment portfolio state. The entire process is rigorously tracked, ensuring complete traceability and consistency throughout the decision-making pipeline. We provide more details on how the workflow effectively coordinates information exchange in Appendix D.1.
Evaluation Interface. As shown in Figure 1(b), we provide a web-based interface for presenting the trading performance for each LLM. Inspired by the previous work, ChatBot Arena [8] and Open FinLLM Leaderboard [28], the interface is designed to be comprehensive and fine-grained, allowing for in-depth analysis of the trading behavior. Please refer to Appendix D.3 for more details.
@@ 3 Experimental Setting
Financial Data Integration. We integrate upstream data from well-known and trusted financial provider APIs (e.g., Alpha Vantage, Yahoo Finance). The data covers not only granular ticker-level information, such as financial statement, company news, daily trading statistics, and insider transactions, but also the macro indicators and policy news. Parametric settings are presented in Appendix D.4 for more details.
@@
OpenAI GPT-4.1 [40] Apr 2025 June 2024
Meta Llama 4 Scout [38] Apr 2025 Aug 2024
Google Gemini 2.5 Flash [14] Apr 2025 Jan 2025
Anthropic Claude 3.7 Sonnet [2] Feb 2025 Oct 2024
xAI Grok 3 mini Beta [51] Feb 2025 Nov 2024
DeepSeek DeepSeek-V3 [10] Mar 2025 Dec 2024
Alibaba Qwen2.5-Max [1] Jan 2025 NA
ByteDance Doubao-1.5-pro [6] Jan 2025 NA
Zhipu GLM-4-Air [61] Apr 2025 NA
@@
Signal and Decision Validity. The validity of signals and decisions generated by DeepFund was evaluated over 24 trading days. The signal and decision are regarded as valid only if the justification is correctly provided. Failing to do so will result in No signal provided due to error or Just hold due to error, respectively. Out of a total of 4320 signals and 1080 trading decisions, DeepFund successfully produced 4144 signals (96% validity) and 1059 trading decisions (98% validity). Such high validity rates indicate the robustness and reliability in generating timely and actionable outputs. Further detailed statistics are available in Appendix B.
Evaluation Metrics. We employ standard financial metrics to measure performance rigorously, including Cumulative Return ($\mathbf{CR}$[16], Cumulative Return at Buy & Hold ($\mathbf{CR}_{bnh}$[12], Sharpe Ratio ($\mathbf{SR}$[43], Maximum Drawdown ($\mathbf{MDD}$[36], Win Rate ($\mathbf{WR}$[7], Beta ($\beta$[21], and Alpha ($\alpha$[21]. The detailed definitions and formulas for these metrics can be found in Appendix C.
Implementation Details. We build the agentic workflow via LangChain, a powerful toolkit for building LLM-based applications. All LLM inferences occur through provider-specific APIs, using standardized prompts (see Appendix E) and the same temperature for fairness (refer to Appendix D.4 for more details). Additionally, we utilize Supabase, a PostgreSQL-based cloud database, to store all activities, including historical decisions, portfolio states, and analytical signals. The evaluation incurred approximately \$100 in total costs (i.e., LLM APIs 40%, financial data 40%, and cloud database 20%, correspondingly). To support reproducibility, our code repository is publicly available.
@@ 4.4 Trading Personalities Unveiled: Profiling the LLM Investor (Q4)
Figure 5: Composition portfolio value for DeepSeek and Grok during the trading period. It shows the holding value of each ticker in portfolio and the remaining cashflow.
@@ 5 Related Work
Benchmarking LLMs in Financial Domain. With the growing application of LLMs in finance, numerous benchmarks have been developed to evaluate their capabilities in financial contexts. Financial LLM benchmarks have evolved from document-understanding frameworks like TAT-QA [62], FinanceBench [18], CFBenchmark [24], and FinNLI [37] to investment decision-making evaluations such as FinRL-Meta [33], FinBen [53] and InvestorBench [25]. These benchmarks share a common focus on assessing LLMs financial knowledge and reasoning capabilities, but still face a fundamental challenge that temporal mismatch between model pre-training data and evaluation window leads to either information leakage when models are tested on historical data they’ve been trained on, or incomplete evaluation when tested on periods beyond their knowledge boundary. Our work shifts from static evaluation to dynamic evaluation to tackle this challenge.
Live Benchmarking. It has been a growing interest in developing live benchmark for AI systems to eliminate the “time-travel” problem in recent years. Several works have explored contamination-free evaluation approaches in general domains, such as LiveCodeBench [19], ForecastBench [22] and LiveBench [50]. These benchmarks are designed to evaluate LLMs with in-context learning, code generation, and domain-specific tasks, with regular updates to ensure the benchmark is always up-to-date. Particularly in the financial domain, FinRL-Meta [33] builds benchmarks for reinforcement learning approaches, and ForecastBench [22] explores the evaluation of LLMs forecasting capabilities through question-answering that covers market-related questions. With the advancement of multi-agent systems, several works [48, 45] have developed a self-evolving multi-agent framework for dynamic evaluation on LLMs. However, in the specific domain of stock trading and fund investment, existing benchmarks have largely remained static, relying on historical data and back-testing approaches. Our work represents the first work that achieves true live benchmarking for fund investment.
@@ 6 Conclusion
DeepFund is a powerful benchmarking tool for evaluating and comparing the performance of various LLMs in the context of real-time fund investment. It provides a standardized multi-agent trading workflow with a connection to a live environment and LLM factory. Besides, we have conducted empirical studies to show the effectiveness of our framework and reveal the potential of LLMs in fund investment. Overall, DeepFund creates a new paradigm for evaluating LLMs in fund investment, which could contribute to the development of reliable and effective financial AI tools.
@@ 7 Limitation
The current implementation highly simplifies the trading process and does not account for many practical considerations such as transaction fees, market trading restrictions, and hybrid trading strategies. These details could potentially impact trading performance, but have not been evaluated in our current framework. Besides, we currently only focus on the US stock market. Future iterations should expand to include a more diverse range of markets, such as non-US stocks, derivatives, commodities, and alternative investments to better assess LLMs’ performance across different assets. Lastly, the evaluation period was short and occurred during a declining market, which could skew our results toward specific trading approaches. Conducting longer tests across various market conditions (e.g., rising, falling, and stable markets) would give more reliable and applicable insights into LLMs’ investment abilities.
@@ A.2 The Cost Efficiency on OpenAI Family
Appendix B Signal and Decision Statistics

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated notable capabilities across financial tasks, including financial report summarization, earnings call transcript analysis, and asset classification. However, their real-world effectiveness in managing complex fund investment remains inadequately assessed. A fundamental limitation of existing benchmarks for evaluating LLM-driven trading strategies is their reliance on historical back-testing, inadvertently enabling LLMs to “time travel” – leveraging future information embedded in their training corpora, thus resulting in possible information leakage and overly optimistic performance estimates. To address this issue, we introduce DeepFund, a live fund benchmark tool designed to rigorously evaluate LLM in real-time market conditions. Utilizing a multi-agent architecture, DeepFund connects directly with real-time stock market data – specifically data published after each model’s pretraining cutoff – to ensure fair and leakage-free evaluations. Empirical tests on nine flagship LLMs from leading global institutions across multiple investment dimensions—including ticker-level analysis, investment decision-making, portfolio management, and risk control—reveal significant practical challenges. Notably, even cutting-edge models such as DeepSeek-V3 and Claude-3.7-Sonnet incur net trading losses within DeepFund real-time evaluation environment, underscoring the present limitations of LLMs for active fund management. Our code is available at https://github.com/HKUSTDial/DeepFund.
@@ 1 Introduction
The financial industry has witnessed an AI-driven revolution over the past decade [31, 36, 69, 58]. Advanced AI techniques, particularly Large Language Models (LLMs), have transformed practices across multiple domains, including high-frequency trading algorithms [41, 5, 76, 37], risk assessment models [13, 70, 21], investment decisions [63, 32, 66, 68, 30], and data analysis [56, 26, 46, 39, 47, 28], thereby fundamentally reshaping how financial institutions operate and make decisions.
@@
A fundamental limitation in extending these evaluations to trading performance lies in their reliance on retrospective back-testing. Back-testing is the standard method for assessing trading strategies [3, 16], but it becomes problematic when applied to LLM-driven strategies because the model may have been pre-trained on the very historical data used for testing, which leads to a severe information leakage issue [65, 54, 51, 11]. Undoubtedly, an LLM can appear to perform extraordinarily well on historical market data simply by regurgitating events it has already seen, rather than genuinely predicting outcomes [42, 14, 73, 74]. In other words, the model can effectively “time travel” by using future knowledge during evaluation—a form of cheating that inflates its apparent performance. This issue is exacerbated by the varying knowledge cut-off dates of different LLMs.
@@
DeepFund: Fund Benchmark Going Live. To address the critical gap identified above, we introduce DeepFund – a comprehensive framework for real-time fund investment benchmarking, as shown in Figure 1(b). Inspired by previous works [38, 20, 61, 75, 45], DeepFund assesses LLM’s ability to make effective investment decisions in a live-market environment, explicitly preventing any leakage of future data. In particular, our approach offers three key contributions:
@@ 2 DeepFund: Multi-Agent Fund Investment Going Live
To facilitate seamless and flexible data ingestion, we implement a modular API gateway that interfaces with multiple financial data providers, such as Yahoo Finance and Alpha Vantage. This modularity (refer to Appendix B.5) ensures adaptability to diverse data sources and straightforward integration into varying market contexts. By offering direct feedback from live market conditions, the environment guarantees authenticity in the decision-making process, fundamentally shifting the evaluation paradigm from retrospective back-testing to dynamic, real-time interaction.
Single Agent Design. Similar to previous works [32, 63, 30, 43, 40], each agent in our framework, powered by the selected LLM backend, fulfills a specific role within the investment process:
@@
Portfolio Manager: Integrates multiple analyst signals to make executive investment decisions (Buy, Sell, Hold), manages risk control (i.e., the portion of holdings and cash), and maintains a dual-memory architecture (see Appendix B.2) to reflect historical transactions and current portfolio states.
@@
Multi-Agent Workflow. The multi-agent workflow [67, 35, 60, 44, 27] adopts an orchestrator-worker paradigm [^1] See Anthropic blog https://www.anthropic.com/engineering/building-effective-agents to mimic a realistic fund management process. Initially, the Financial Planner selects and assigns analysts based on real-time market conditions and portfolio status. Next, the selected Analyst Team concurrently processes domain-specific information and generates structured analytical signals. Finally, the Portfolio Manager synthesizes these signals, evaluates portfolio risks, decides optimal trading actions, and updates the investment portfolio state. The entire process is rigorously tracked, ensuring complete traceability and consistency throughout the decision-making pipeline. We provide more details on how the workflow effectively coordinates information exchange in Appendix B.1.
Evaluation Interface. As shown in Figure 1(b), we provide a web-based interface for presenting the trading performance for each LLM. Inspired by the previous work, ChatBot Arena [8] and Open FinLLM Leaderboard [33], the interface is designed to be comprehensive and fine-grained, allowing for in-depth analysis of the trading behavior. Please refer to Appendix B.3 for more details.
@@ 3 Experimental Setting
Financial Data Integration. We integrate upstream data from well-known and trusted financial provider APIs (e.g., Alpha Vantage, Yahoo Finance). The data covers not only granular ticker-level information, such as financial statement, company news, daily trading statistics, and insider transactions, but also the macro indicators and policy news. Parametric settings are presented in Appendix B.4 for more details.
@@
OpenAI GPT-4.1 [52]
Apr 2025 June 2024
Meta Llama 4 Scout [50]
Apr 2025 Aug 2024
Google Gemini 2.5 Flash [15]
Apr 2025 Jan 2025
Anthropic Claude 3.7 Sonnet [2]
Feb 2025 Oct 2024
xAI Grok 3 mini Beta [62]
Feb 2025 Nov 2024
DeepSeek DeepSeek-V3 [10]
Mar 2025 Dec 2024
Alibaba Qwen2.5-Max [1]
Jan 2025 NA
ByteDance Doubao-1.5-pro [6]
Jan 2025 NA
Zhipu GLM-4-Air [71]
Apr 2025 NA
@@
Signal and Decision Validity. The validity of signals and decisions generated by DeepFund was evaluated over 24 trading days. The signal and decision are regarded as valid only if the justification is correctly provided. Failing to do so will result in No signal provided due to error or Just hold due to error, respectively. Out of a total of 4320 signals and 1080 trading decisions, DeepFund successfully produced 4144 signals (96% validity) and 1059 trading decisions (98% validity). Such high validity rates indicate the robustness and reliability in generating timely and actionable outputs. Further detailed statistics are available in Appendix C.
Evaluation Metrics. We employ standard financial metrics to measure performance rigorously, including Cumulative Return ($\mathbf{CR}$) [17], Cumulative Return at Buy & Hold ($\mathbf{CR}_{bnh}$) [12], Sharpe Ratio ($\mathbf{SR}$) [55], Maximum Drawdown ($\mathbf{MDD}$) [48], Win Rate ($\mathbf{WR}$) [7], Beta ($\beta$) [22], and Alpha ($\alpha$) [22]. The detailed definitions and formulas for these metrics can be found in Appendix D.
Implementation Details. We build the agentic workflow via LangChain, a powerful toolkit for building LLM-based applications. All LLM inferences occur through provider-specific APIs, using standardized prompts (see Appendix E) and the same temperature for fairness (refer to Appendix B.4 for more details). Additionally, we utilize Supabase, a PostgreSQL-based cloud database, to store all activities, including historical decisions, portfolio states, and analytical signals. The evaluation incurred approximately \$100 in total costs (i.e., LLM APIs 40%, financial data 40%, and cloud database 20%, correspondingly). To support reproducibility, our code repository is publicly available.
@@ 4.4 Trading Personalities Unveiled: Profiling the LLM Investor (Q4)
Figure 5: Composition portfolio value for DeepSeek and Grok during the trading period. It shows the holding value of each ticker in the portfolio and the remaining cash flow.
@@ 5 Related Work
Benchmarking LLMs in Financial Domain. With the blossoming research on LLMs in finance [63, 32, 30], numerous benchmarks have been developed to evaluate their capabilities in financial contexts. Financial LLM benchmarks have evolved from document-understanding frameworks like TAT-QA [72], FinanceBench [19], CFBenchmark [25], and FinNLI [49] to investment decision-making evaluations such as FinRL-Meta [38], FinBen [64] and InvestorBench [29]. These benchmarks share a common focus on assessing LLMs’ financial knowledge and reasoning capabilities, but still face a fundamental challenge that the temporal mismatch between model pre-training data and the evaluation window leads to either information leakage when models are tested on historical data they’ve been trained on, or incomplete evaluation when tested on periods beyond their knowledge boundary. Our work shifts from static evaluation to dynamic evaluation to tackle this challenge.
Live Benchmarking. There has been a growing interest in developing live benchmarks for AI systems to eliminate the “time-travel” problem in recent years. Several works have explored contamination-free evaluation approaches in general domains, such as LiveCodeBench [20], ForecastBench [23], and LiveBench [61]. These benchmarks are designed to evaluate LLMs with in-context learning, code generation, and domain-specific tasks, with regular updates to ensure the benchmark is always up-to-date. Particularly in the financial domain, FinRL-Meta [38] builds benchmarks for reinforcement learning approaches, and ForecastBench [23] explores the evaluation of LLMs’ forecasting capabilities through question-answering that covers market-related questions. With the advancement of multi-agent systems, several works [59, 57] have developed a self-evolving multi-agent framework for dynamic evaluation on LLMs. However, in the specific domain of stock trading and fund investment, existing benchmarks have largely remained static, relying on historical data and back-testing approaches. Our work represents the first work that achieves true live benchmarking for fund investment.
@@ 6 Conclusion
DeepFund is a novel benchmarking tool for evaluating and comparing the performance of various LLMs in the context of real-time fund investment. It provides a standardized multi-agent trading workflow with a connection to a live environment and LLM factory. Besides, we have conducted empirical studies to show the effectiveness of our framework and reveal the notable potential of LLMs in fund investment. Overall, DeepFund creates a new paradigm for evaluating LLMs in fund investment, which could contribute to the development of reliable and effective financial AI tools.
@@ 7 Limitation
The current implementation highly simplifies the trading context (i.e., US stock market only) and does not account for many practical considerations such as transaction fees, market trading restrictions, and hybrid trading strategies. These details could potentially impact trading performance, but have not been evaluated in our current framework. Therefore, we aim to incorporate index-aligned universes, realistic execution frictions, and broader market conditions in future iterations. Meanwhile, the analysis depth can be further improved by adding more index-based (e.g., relative performance to market indices) and LLM-based (e.g., reasoning cost, consistency, explainability) evaluation metrics. Lastly, the evaluation period was short and occurred during a volatile market, which could skew our results toward specific trading approaches. Conducting longer tests across various circumstances (e.g., bullish, bearish and volatile markets) would manifest more reliable and applicable insights into LLMs’ investment abilities.
8 Broader Impacts
9 Ethical Statement
10 Acknowledgements
This paper was supported by Young Talent Support Project of Guangzhou Association for Science and Technology (QT-2025-001); the NSF of China (62402409); Guangdong Basic and Applied Basic Research Foundation (2023A1515110545); Guangzhou Basic and Applied Basic Research Foundation (2025A04J3935); Guangzhou-HKUST(GZ) Joint Funding Program (2025A03J3714); and Guangdong Provincial Project (2023CX10X008). Lastly, we genuinely thank the reviewers for their valuable comments and suggestions during the rebuttal period.
@@ A.2 The Cost Efficiency on OpenAI Family
A.3 Extended Trading Performance to entire Q2 2025
To address the concern on the narrow evaluation window, we expanded our live trading window to cover the entire Q2 2025 (see Table 4). Each LLM continues to actively manage its portfolio and surf in the market, with Grok-3 still maintaining its leading profitability. Some LLMs such as GPT 4.1, Claude 3.7, and DeepSeek V3 achieved net profits eventually, while other LLMs such as GLM 4, Qwen Max, and Gemini 2.5 still incur losses. Results from such a longer period remain consistent with our original findings, which further supports the robustness of our conclusions.
Table 4: Averaged Weekly Asset Total Value by LLM in Q2 2025. Dates are the end of each week.
Date GPT 4.1 Claude 3.7 Grok 3 Llama 4 Gemini 2.5 DeepSeek V3 Qwen Max Doubao 1.5 GLM 4
2025-04-06 100207.56 99370.47 101031.59 100200.65 99819.65 99305.63 99973.54 99387.25 99629.81
2025-04-13 92824.93 93933.56 98978.32 94181.13 96280.95 91682.35 92652.86 91204.98 91695.78
Below are new results
2025-04-20 94322.82 96202.42 101156.52 95466.44 98212.7 94351.44 93099.94 92098.54 92523.98
2025-04-27 94993.84 97186.15 101527.47 96768.25 98205.38 96133.87 94084.57 93030.82 93893.19
2025-05-04 96627.35 98242.37 102151.96 98289.12 98349.17 97987.88 94765.69 94182.54 95209.34
2025-05-11 96247.15 98319.67 102227.45 97770.66 97374.24 98223.28 94376.73 93827.1 95528.87
2025-05-18 98815.25 100726.09 103425.42 102377.7 98323.0 101482.47 95853.75 97527.84 98038.4
2025-05-25 97701.93 99340.52 103385.97 100680.71 97822.53 99767.74 94142.15 96939.64 96392.79
2025-06-01 98589.52 99419.7 103625.78 100703.5 98473.83 99969.97 94540.49 97865.15 96501.95
2025-06-08 99292.85 99914.3 103972.52 101252.86 98746.37 100703.84 94868.62 98584.24 96893.33
2025-06-15 99942.06 100360.74 104263.5 101935.87 99192.8 101416.86 95380.01 99262.71 97060.48
2025-06-22 99470.03 99668.47 104700.82 101698.64 98919.57 100873.6 95216.37 98985.76 96356.1
2025-06-29 101442.76 101234.08 106262.44 102859.48 99762.68 102851.0 96383.33 100865.84 97670.72
Appendix B Technical Details
B.1 Data Models and Operational Schemas
B.2 Memory Management
B.3 Evaluation Interface
The interface serves as the first AI Live Investment Arena, designed to evaluate the trading and investment capabilities of various Large Language Models (LLMs) across financial markets. It provides a comprehensive environment for assessing how LLMs can ingest financial information, drive multi-agent systems, and make trading decisions in real-world market scenarios. You can explore its features further by visiting https://deepfund.paradoox.ai/. The main features are:
Table 5: Key features of the DeepFund valuation interface.
Feature Name Description
B.4 Parameter Settings
Table 6: DeepFund parameter settings.
Parameter Default Value Usage
LLM temperature 0.5 Control the randomness of the LLM inference.
Retry times 3 Number of retries for LLM inference.
Technical window 100 OHLCV covered days for technical analysis.
Insider count 10 Insider transactions for insider analysis.
Number of news 10 Compnany news and policy analysis.
B.5 Systematic Scalability
Appendix C Signal and Decision Statistics